# Restyling Unsupervised Concept Based Interpretable Networks with Generative Models

### v1

@@ 5 Conclusion
We introduced a novel architecture for Visualizable CoIN (VisCoIN), that addresses major limitations to visualize unsupervised concept dictionaries learnt in CoIN systems for large-scale images. Our architecture integrates the visualization process in the pipeline of the model training, by leveraging a pretrained generative model using a concept translator module. This module maps concept representation to the latent space of the fixed generative model. During training, we additionally enforce a viewability property that promotes reconstruction of high-quality images through the generative model. Finally, we defined new evaluation metrics for this novel interpretation pipeline, to better align evaluation of concept dictionaries and interpretations provided to a user. Future works include adapting the design of this system for other type of generative models, or applying it to different visual domains.
Appendix 0.A Design choices considered with the generative model
0.A.1 Desiderata for generative model
We discuss below the reasoning and requirements governing our choice of generative model $G$, i.e., a pretrained StyleGAN2-ADA.
  • The generative model should have the ability to model the input with a low dimensional latent space (compared to input dimensions), since a concept representation is typically much lower dimensional than input. Even so, it must possess a structured latent space so that modifying a concept activation can enable latent traversal for visualization. These requirements discourage the use of invertible neural networks [10, 18], or diffusion models [64, 57] whose structure of latent spaces is still being explored [42, 53]. Furthermore, since we constantly rely on input reconstruction during training, the slow sampling process in diffusion models for large-scale images also hinders their suitability for our use case.
@@ 0.A.1 Desiderata for generative model
  • One crucial role of the generative model is to generate high quality natural images. This is essential not only for better reconstruction of large-scale images, but also crucial to ground any visual modifications in the generated images back to the original input. This encouraged our preference for GANs instead of Variational Auto-Encoders (VAEs) [39, 54], specially for large-scale images.
  • The generative model also needs to be able to train in limited data regimes if a pretrained generator is not readily available.
Hence, we opted for a competitive, flexible and widely used generative model, StyleGAN2-ADA [31], shown to have demonstrated high-quality generation for various image domains.
0.A.2 Defining concepts in stylespace of pretrained StyleGAN
Previous works using StyleGAN for interpretation deliberately choose to use the Stylespace as concept dictionary [43], which is considered the most disentangled and interpretable latent space of a StyleGAN [72]. However, in regard to a by-design interpretable architecture, there is a crucial issue in using the Stylespace of a pretrained StyleGAN as the concept feature space. Indeed, the size of Stylespace, even for small-scale images, runs into many thousands, which is much larger than the size of typical dictionaries in CoINs (by a factor of 10 to a 100). This directly compromises the interpretable structure of the prediction model [46] with much lower number of classes for prediction. Moreover, since the model is pretrained for generation, the space doesn’t have sparse activations, further worsening the issue. Note that with the same reasoning, we considered using the extended latent space $\mathcal{W}^{+}$ as concept dictionary unsuitable. The final option that remains is using $\mathcal{W}$ space as our concept dictionary. While at first glance it seems an interesting option, since its size is comparable to concept dictionaries in CoINs, and it is also considered as better structured than $\mathcal{W}^{+}$ for latent traversals [73], inverting images to $\mathcal{W}$ space has been found to be lot more challenging [1]. Doing so under even more constraints such as those of concept dictionaries in CoIN, poses an even harder challenge. Hence, we adhered to the CoIN structure to learn a small dictionary compared to the size of $\mathcal{W}^{+}$. Given the constraints imposed on $\Phi$, we designed the decoding part to prioritize better reconstruction.
Appendix 0.B Further system details
0.B.1 Network architectures
@@
We already discuss in the main text the architectures for $f$ and $G$, both using standard and widely used models, ResNet50 [26] and StyleGAN2-ADA [31] respectively. $f$ and $G$ are pretrained and fixed during training of VisCoIN. As part of our training we train three subnetworks $\Psi,\Omega,\Theta$. We already described $\Theta$ in the main text, as consisting of a pooling (maxpool), linear and softmax layers in the respective order. We describe the architecture of $\Psi$ network that operates on hidden layers of $f$ and computes the dictionary activations $\Phi(x)$ and the supporting representation $\Phi^{\prime}(x)$ and the design of concept translator $\Omega$ that maps $(\Phi(x),\Phi^{\prime}(x))$ to extended latent space $\mathcal{W}^{+}$ of StyleGAN to reconstruct $x$.
0.B.1.1 Architecture of $\Psi$
@@
Our architecture of $\Psi$ mostly follows proposed architecture of $\Psi$ for FLINT [51] which accesses output of two layers for ResNet18 close to the output layer (output of block 3 and penultimate layer of block 4). The ResNet50 also follows a similar structure with 4 blocks. Each block however contains 3, 4, 3 and 3 sub-blocks termed “bottleneck” respectively. In terms of the set of layers accessed by $\Psi$ for VisCoIN, in addition to the corresponding two layers in ResNet50 (output of block 3 of shape $1024\times 16\times 16$, output of penultimate bottleneck layer in block 4 of shape $2048\times 8\times 8$), we also access a third layer for improved reconstruction (output of block 2 of shape $512\times 32\times 32$). Each layer output is passed through a convolutional layer and brought to a common shape of $512\times 8\times 8$, the lowest resolution and feature maps. We then concatenate all the feature maps and create two branches, one that outputs $\Phi(x)$ and the other outputs $\Phi^{\prime}(x)$. The branch computing $\Phi(x)$ is simply two convolutional and a pooling layer yielding an output shape of $K\times 3\times 3$, where $K$ is the number of concepts, and each $\phi_{k}(x)$ is a convolutional map of size $3\times 3$. Thus, the total number of elements in each $\phi_{k}(x)$ is $b=9$. The branch computing support representation $\Phi^{\prime}(x)$ (unconstrained except by reconstruction) applies two fully connected layers to output same number of elements as in $\Phi(x)$. The usefulness of the support representation $\Phi^{\prime}(x)$ is analysed in Tab. 7.
0.B.1.2 Architecture of $\Omega$
@@
For translating the concept activations $\Phi(x)$ and the supporting representation $\Phi^{\prime}(x)$ to the extended latent space, we use the concept translator $\Omega$. As described in the main text, it consists of single fully-connected layers, one for each latent vector (14 different vectors of size 512 for resolution $256\times 256$). We control prediction of a small subset of latent vectors using $\Phi^{\prime}$, which are expected to contain minimal classification information. To design which latent vectors to control using $\Phi^{\prime}$, we relied on findings from the work in [35], which roughly divides the different latent vectors in $\mathcal{W}^{+}$ as controlling the coarse, mid and fine level features of the generated image with increasing resolution. The first 4 latent vectors control typically the coarse details, such as view-point, pose or scale. The next 4 roughly control the mid-level details, such as shape or background, while the rest control the fine-level details, such as textures and colours. We expect the relevant features for classification to be controlled mostly by mid-level and fine-level latent vectors, except possibly for the highest resolution where very fine-scaled details are controlled. In particular, we predict the first three and last two style vectors using $\Phi^{\prime}(x)$, which are expected to contain less relevant information for our concept representation. The rest of the style vectors (9 out of 14 for resolution 256) are predicted from $\Phi(x)$. Note that most of the style vectors are still predicted using $\Phi(x)$ to preserve its importance in reconstruction. Also note that employing single FC layers associates each concept activation with a direction in extended latent space.
0.B.2 Training details
@@
The steps to train our system on a given dataset can be divided into three modular parts: (1) Obtaining a pretrained classifier $f$ with “strong” performance that can provide high-quality source representations to learn from, (2) Obtaining a pretrained generator $G$ that can approximate well the distribution of the given dataset, and (3) Training of $g$ with VisCoIN using the pretrained $f$ and $G$. When a pretrained $f$ or $G$ is not easily available, we train them on their respective tasks on the given dataset. Among our 3 datasets, CelebA-HQ [30], CUB-200 [70] and Stanford Cars [41], we easily found a pretrained $G$ for CelebA-HQ. All other combinations of $f$ and $G$ were pretrained. We describe the training details of $f$, $G$ and VisCoIN below:
0.B.2.1 Pretraining $f$
@@
We pretrain $f$ for classification on each of our datasets before using it for training VisCoIN. We use Adam optimizer [38] with fixed learning rate 0.0001 on CUB-200 and 0.001 on CelebA-HQ to train $f$. On Stanford-Cars, we use SGD optimizer with a starting learning rate of 0.1, decayed by a factor 0.1 after 30 and 60 epochs. The training is initialized with pretrained weights from ImageNet in each case, and fine-tuned for 10, 30 and 90 epochs on CelebA-HQ, CUB-200 and Stanford-Cars respectively. In all cases, during pretraining, the images are resized to size $256\times 256$.The accuracy of $f$ is already reported in the main paper. All of these experiments have been conducted on a single A100 GPU, with a batch size of 64 for CelebA-HQ and 128 for Stanford-Cars dataset and on V100 GPU with a batch size of 32 for CUB-200.
0.B.2.2 Pretraining $G$
@@
We use a pretrained StyleGAN2-ADA [31] for all our experiments. On CelebA-HQ, we used a pretrained checkpoint available from NVIDIA. For CUB-200 and Stanford-Cars, we pretrain $G$ ourselves. Note that since we want $G$ to generate images entirely from information provided by $\Phi(x)$, we do not use any class labels when training $G$. One can face two separate challenges in these limited data regimes. The first is to train $G$ in a stable fashion. Our solution to this, as discussed previously, comes from using the ADA training strategy for StyleGAN2. We also use the official StyleGAN2-ADA Pytorch repository to train our models.https://github.com/NVlabs/stylegan2-ada-pytorch A second challenge can come from limitations to training resources since these models might require to be trained with tens of millions of real/dataset images (“shown” to the discriminator) in order to reach high quality generation. This could potentially require training with multiple GPUs for multiple days. We address this issue to a reasonable extent by fine-tuning pretrained checkpoints. We utilize the insights from [24] and fine-tune a checkpoint from ImageNet for CUB-200, and LSUN Cars [79] for Stanford Cars. The choices of these specific models was specifically based on the idea that these datasets were the closest domains we had access of pretrained checkpoints to.
@@
0.B.2.3 Details for VisCoIN
We train for 50K iterations on CelebA-HQ and 100K iterations on CUB-200 and Stanford-Cars. We use Adam optimizer with learning rate $0.0001$ for all subnetworks and on all datasets. During training, each batch consists of 8 samples from the training data and 8 synthetic samples randomly generated using $G$. This practice of utilizing the synthetic samples from $G$ is fairly common for encoder-based GAN inversion systems [75, 73], and an additional advantage for our system to use a pretrained $G$. Note that the use of fidelity loss with a pretrained $f$ instead of a classification loss on $g(x)$ fits neatly with this, as one cannot obtain any ground-truth annotations for the synthetic samples. The training data samples use a random cropping and random horizontal flip augmentation in all cases. All images are normalized to the range $[-1,1]$ and have resolution $256\times 256$ for processing. This is the default range and resolution we use for pretraining for $f$ and $G$ too. We have already described the architectures of all our components, pretrained or trained as part of training VisCoIN. We tabulate below in Tab. 3 the hyperparameter values for all our datasets. To limit the amount of hyperparameters to tune, we used a fixed $\alpha=0.5$ (weight for output fidelity loss) and $\beta=3$ (weight for LPIPS loss) for all datasets. The rationale behind choice of all hyperparameters is discussed in Appendix 0.D, wherein we also present the ablation studies w.r.t to multiple components.

### v2

@@ 5 Conclusion
We introduced a novel architecture for Visualizable CoIN (VisCoIN), that addresses major limitations to visualize unsupervised concept dictionaries learnt in CoIN systems for large-scale images. Our architecture integrates the visualization process in the pipeline of the model training, by leveraging a pretrained generative model using a concept translator module. This module maps concept representation to the latent space of the fixed generative model. During training, we additionally enforce a viewability property that promotes reconstruction of high-quality images through the generative model. Finally, we defined new evaluation metrics for this novel interpretation pipeline, to better align evaluation of concept dictionaries and interpretations provided to a user. Future works include adapting the design of this system for supervised CoINs, multimodal generative models, or extending its application to different data modalities.
Reproducibility Statement
Throughout the paper, we made sure that all our experiments were fully reproducible, describing in details all datasets and architectures considered in Section 4. We then explain the design of $\Psi$ and $\Omega$, training settings, hyperparameters and computation of evaluation metrics in Appendix C.
Acknowledgements
This work was supported by the ANR LIMPID and ANR Far-See. The authors would also like to thank Thibaut de Saivre and Hugo Aoyagi for restructuring the codebase.
Appendix A Language-based concepts
A.1 Supervised CoINs and language model based CBMs
Recent variants of concept bottleneck models (CBMs) are trained with concept sets generated from language models (Yuksekgonul et al., 2022; Oikarinen et al., 2023; Yang et al., 2023b; Panousis et al., 2024). These models are also CoINs that do not require human annotations for training. They obtain sample-wise concept annotations using concept set constructed from language models. Specifically, for each class, they prompt a language model to generate a set of text descriptions describing the class. This set of descriptions is then used as the concept bank. For each sample, the concept annotation during training is obtained by CLIP similarity between the image and text descriptions of its ground-truth class. Note that all of these steps are executed prior to training the concept bottleneck model.
Interestingly, while these recent CBMs provide text descriptions for concepts thanks to automatic extraction from LLMs, they are even worse offenders of the visualization limitations highlighted in Section 1. This is because they currently do not have any reliable visualization pipeline to confirm if the visual detection of a given concept by the underlying CBM corresponds to its text description or not. Moreover, in many cases the concepts as part of concept annotations are not even visually grounded (e.g. “Loyal/honest” for class “Dog”) in which case it is impossible to visualize the concept in the first place. However, they are methodologically very close to original CBMs or supervised CoINs as they still use sample-wise annotations to train $\Phi(x)$.
Besides the use of concept labels, CBMs do not have a decoder model. There is no way to conduct any of the main evaluation (reconstruction, faithfulness, consistency) without a decoder, since we need to approximate the input from concept activations to do any of these. All unsupervised CoINs have a decoder and none of the CBMs do, as a result of their training methodological differences. Thus, they are currently out of scope for VisCoIN which is focused more on unsupervised CoINs.
A.2 On the usefulness of visualization
We present below arguments why visualizing concepts is still important:
  • When considering expert or domain specific datasets like Stanford Cars (Krause et al., 2013) (for car models classification) or the MVTec Anomaly Detection (Bergmann et al., 2021) (for anomaly detection of object in production lines) for instance, visualizing concepts directly on the objects is simpler and faster to understand for human operators, rather than reading a text description.
  • For certain computer vision applications (eg. self-driving cars, medical imaging tasks), visualization provides spatially localized interpretations, which is more difficult and cumbersome with text. For instance, if a concept relating to “red light” is activated for an image, to get a thorough understanding of the model’s decision, it is crucial to identify which regions and what content in the image activates the concept. Ideally, this would be best proved by a human evaluation comparing language descriptions and visualizations for real-world applications, to evaluate if/how much advantage visualizations add. However, for fair comparison language description and visualization should be of the same concept dictionary for the same model. To design such a study and system remains a challenging problem, which we will explore as a future direction.
  • The LLMs/VLMs which the recent CBMs are based on (particularly CLIP (Radford et al., 2021)) are limited when detecting concepts and image details at a finer spatial scale (Gou et al., 2024).
  • As discussed above, the current methods are prone to generating concept descriptions not grounded in any visual information, which also harms their interpretability.
  • In the case of LLM/VLM based CBMs, there are also concerns about faithfulness of concept detection to the text description. This is a similar issue to concept leakage (Havasi et al., 2022). We believe that the ideas presented in our work, such as viewability, can help in identifying such issues in LLM/VLM based CBMs.
Appendix B Desiderata for generative model
We discuss below the desired properties that can influence the choice of an appropriate generative model $G$, as previously mentioned in Section 3.2.
  1. (i)
    The generative model should have the ability to model the input with a low dimensional latent space (compared to input dimensions), since a concept representation is typically much lower dimensional than input.
  2. (ii)
    It possesses a structured latent space that admits convenient mechanisms for meaningful latent traversal. This directly helps in designing a $\Omega$ such that modifying a concept activation can enable latent traversal for visualization.
  3. (iii)
@@ Appendix B Desiderata for generative model
    The generative model $G$ is able to unconditionally generate high quality samples, close to the underlying dataset/data distribution. This is essential not only for better reconstruction of images, but also crucial to ground any visual modifications in the generated images back to the original input.
The desiderata discussed above enables our approach to be compatible with a variety of generative models. Notably, most GANs and VAEs satisfy the first two requirements. Provided they are capable to approximate well the underlying data distribution, our approach fits well with them as choice of $G$. The experiments in main text focus on StyleGAN2-ADA (Karras et al., 2020) as our $G$ architecture given its high-quality generation for various large-scale image domains. We also illustrate applicability of our method of other $G$ architectures such as ProgressiveGAN (Karras et al., 2018) and $\beta$-VAE (Higgins et al., 2017) in Appendix D.
For diffusion model (Sohl-Dickstein et al., 2015) architectures, despite the recent positive steps towards understanding their latent space (Kwon et al., 2023; Park et al., 2023), it is currently difficult to design an $\Omega$ that allows a straightforward and meaningful latent traversal. Nevertheless, it is worth mentioning that latent diffusion models (Rombach et al., 2022) do satisfy (i) and (iii) in many cases. With further research in understanding their latent spaces, we believe it could be possible to satisfy (ii) and extend VisCoIN to them.
Appendix C Further system details
C.1 Network architectures
@@
The networks $f$ and $G$ are pretrained and fixed during training of VisCoIN. As part of our training, we train three subnetworks $\Psi,\Omega,\Theta$. We already described $\Theta$ in the main text, as consisting of a pooling (maxpool), linear and softmax layers in the respective order.
General Architecture of $\Psi$
@@
Our architecture of $\Psi$ mostly follows proposed architecture of $\Psi$ for FLINT (Parekh et al., 2021) which accesses output of two layers for ResNet18 close to the output layer (output of block 3 and penultimate layer of block 4). The ResNet50 also follows a similar structure with 4 blocks. Each block however contains 3, 4, 3 and 3 sub-blocks termed “bottleneck” respectively. In terms of the set of layers accessed by $\Psi$ for VisCoIN, in addition to the corresponding two layers in ResNet50 (output of block 3 of shape $1024\times 16\times 16$, output of penultimate bottleneck layer in block 4 of shape $2048\times 8\times 8$), we also access a third layer for improved reconstruction (output of block 2 of shape $512\times 32\times 32$). Each layer output is passed through a convolutional layer and brought to a common shape of $512\times 8\times 8$, the lowest resolution and feature maps. We then concatenate all the feature maps to output $\Phi(x)$. We apply two convolutional and a pooling layer yielding an output shape of $K\times 3\times 3$, where $K$ is the number of concepts, and each $\phi_{k}(x)$ is a convolutional map of size $3\times 3$. Thus, the total number of elements in each $\phi_{k}(x)$ is $b=9$.
General Architecture of $\Omega$
The typical theme we use to implement $\Omega$ is as a single linear layer that takes as input $\Phi(x)$ and outputs a vector in the latent space of the generative model. This directly associates each $\phi_{k}$ with a vector in the latent space of $G$, specified by the $k$-th column vector of weight matrix in $\Omega$. While this exactly corresponds to our implementation for ProgressiveGAN, $\beta-$VAE, our network designs for StyleGAN2 have slight modifications to $\Psi,\Omega$ even though we still follow the same themes. We discuss next the precise architectures with StyleGAN next.
C.1.1 Reconstruction with StyleGAN
The reconstruction architecture with StyleGAN is similar to encoder based GAN-inversion architectures used for StyleGAN. Following previous works in this regard (Abdal et al., 2019; Yao et al., 2022), we use the extended latent space $\mathcal{W}^{+}$ for inversion, which corresponds to different latent vectors for different resolutions. We learn the concept translator $\Omega$ to map the concept activations to $\mathcal{W}^{+}$ and bias its computation with average latent vector $\bar{w}$ of $G$.
Since $\Phi(x)$ is a much lower dimensional representation compared to elements in $\mathcal{W}^{+}$ and constrained by losses unrelated to reconstruction, we found it challenging to achieve reconstruction quality close to GAN inversion methods. This issue in principle cannot be completely eliminated without compromising the interpretable predictive structure. However, we alleviate it by learning an unconstrained “supporting” representation as a secondary output from $\Psi$, termed $\Phi^{\prime}(x)$. The only goal for $\Phi^{\prime}(x)$ is to assist $\Omega$ in embedding the input in $\mathcal{W}^{+}$. To predict $\Phi^{\prime}(x)$ the concatenated feature maps in $\Psi$ are sent in a second parallel branch, that applies two fully connected layers to output same number of elements as in $\Phi(x)$. The computation for reconstruction $\tilde{x}$ is then given as:
@@ C.1.1 Reconstruction with StyleGAN
In this case, the concept translator $\Omega$ consists of a set of single fully-connected (FC) layers, one for predicting each latent vector. Each FC layer either takes $\Phi(x)$ or $\Phi^{\prime}(x)$ as input depending upon the latent vector it predicts. To determine, which latent vectors should be controlled by $\Phi^{\prime}$, we rely on findings from the work in Katzir et al. (2022) which roughly divides the different style vectors in $\mathcal{W}^{+}$ as controlling the coarse, mid and fine level features of the generated image with increasing resolution. For 14 latent vectors in case of 256 $\times$ 256 output resolution, it corresponds to 4, 4, 6 vectors respectively. We expect the relevant features for classification to be controlled mostly by mid-level and fine-level style vectors, except possibly for the highest resolution where very fine-scaled details are controlled. Thus, we predict the first three and last two style vectors using $\Phi^{\prime}(x)$. The rest of the style vectors (9 out of 14 for resolution 256) are predicted from $\Phi(x)$. The choice of using $\Phi^{\prime}$ is analyzed with an ablation study in Table 15.
C.2 Training details
@@
The steps to train our system on a given dataset can be divided into three modular parts: (1) Obtaining a pretrained classifier $f$ with “strong” performance that can provide high-quality source representations to learn from, (2) Obtaining a pretrained generator $G$ that can approximate well the distribution of the given dataset, and (3) Training of $g$ with VisCoIN using the pretrained $f$ and $G$. When a pretrained $f$ or $G$ is not easily available, we train them on their respective tasks on the given dataset. Among our 3 datasets, CelebA-HQ (Karras et al., 2018), CUB-200 (Wah et al., 2011) and Stanford Cars (Krause et al., 2013), we easily found a pretrained $G$ for CelebA-HQ. All other combinations of $f$ and $G$ were pretrained. We describe the training details of $f$, $G$ and VisCoIN below.
C.2.1 Pretraining $f$
@@
We pretrain $f$ for classification on each of our datasets before using it for training VisCoIN. We use Adam optimizer (Kingma & Ba, 2014) with fixed learning rate 0.0001 on CUB-200 and 0.001 on CelebA-HQ to train $f$. On Stanford-Cars, we use SGD optimizer with a starting learning rate of 0.1, decayed by a factor 0.1 after 30 and 60 epochs. The training is initialized with pretrained weights from ImageNet in each case, and fine-tuned for 10, 30 and 90 epochs on CelebA-HQ, CUB-200 and Stanford-Cars respectively. In all cases, during pretraining, the images are resized to size $256\times 256$. The accuracy of $f$ is already reported in the main paper. All of these experiments have been conducted on a single A100 GPU, with a batch size of 64 for CelebA-HQ and 128 for Stanford-Cars dataset and on V100 GPU with a batch size of 32 for CUB-200.
C.2.2 Pretraining $G$
@@
We use a pretrained StyleGAN2-ADA (Karras et al., 2020) for experiments in the main text. On CelebA-HQ, we used a pretrained checkpoint available from NVIDIA. For CUB-200 and Stanford-Cars, we pretrain $G$ ourselves. Note that since we want $G$ to generate images entirely from information provided by $\Phi(x)$, we do not use any class labels when training $G$. We use the official StyleGAN2-ADA Pytorch repository to train our models.https://github.com/NVlabs/stylegan2-ada-pytorch A challenge that can arise is from limitations to training resources since these models might require to be trained with tens of millions of real/dataset images (“shown” to the discriminator) in order to reach high quality generation. This could potentially require training with multiple GPUs for multiple days. We address this issue to a reasonable extent by fine-tuning pretrained checkpoints. We utilize the insights from (Grigoryev et al., 2022) and fine-tune a checkpoint from ImageNet for CUB-200, and LSUN Cars (Yu et al., 2015) for Stanford Cars. The choices of these specific models was specifically based on the idea that these datasets were the closest domains we had access of pretrained checkpoints to. We also present experiments with pretrained ProgressiveGAN (Karras et al., 2018) and $\beta$-VAE (Higgins et al., 2017) in Appendix D. For ProgressiveGAN, we use a pretrained checkpoint on CelebA-HQ provided by NVIDIA. The $\beta$-VAE is trained by us with $\beta=2$.
@@
C.2.3 Training VisCoIN
We train for 50K iterations on CelebA-HQ and 100K iterations on CUB-200 and Stanford-Cars. We use Adam optimizer with learning rate $0.0001$ for all subnetworks and on all datasets. During training, each batch consists of 8 samples from the training data and 8 synthetic samples randomly generated using $G$. This practice of utilizing the synthetic samples from $G$ is fairly common for encoder-based GAN inversion systems (Yao et al., 2022; Xia et al., 2022), and an additional advantage for our system to use a pretrained $G$. Note that the use of fidelity loss with a pretrained $f$ instead of a classification loss on $g(x)$ fits neatly with this, as one cannot obtain any ground-truth annotations for the synthetic samples. The training data samples use a random cropping and random horizontal flip augmentation in all cases. All images are normalized to the range $[-1,1]$ and have resolution $256\times 256$ for processing. This is the default range and resolution we use for pretraining for $f$ and $G$ too. We have already described the architectures of all our components, pretrained or trained as part of training VisCoIN. We tabulate below in Table 3 the hyperparameter values for all our datasets. To limit the amount of hyperparameters to tune, we used a fixed $\alpha=0.5$ (weight for output fidelity loss) and $\beta=3$ (weight for LPIPS loss) for all datasets. The rationale behind choice of all hyperparameters is discussed in Appendix F, wherein we also present the ablation studies w.r.t to multiple components.